4) A free proton has a wave function ψ(x)-A sin (kx), where k = 1.2 x 1010 m-1 a. What is the protons λ? b. What is the protons momentum? c. What is the protons speed? d. Normalize v(x) if the wave only exists inside an infinite square well with width a 2.1 m, (so that Ψ(x) = A sin (kx) between 0 < x < a and v(x)-0 otherwise).
Answer #1

a] The wavelength of the proton will be:

\lambda = \frac{2\pi}{k} = \frac{2\pi}{1.2\times 10^{10}} = 5.236 \times 10^{-10}m

b] Proton's momentum is given by:

p = \frac{h}{\lambda}=\frac{6.626\times 10^{-34}}{5.236\times 10^{-10}} = 1.2655\times 10^{-24}kgm/s

c] p = mv

m = 1.67 x 10-27 kg

=> v = 757.768 m/s

d] The Normalization condition here will be:

\int^{2.1}_0 \psi (x)\psi^*(x)dx = 1

so, A\int^{2.1}_0 sin^2(kx)dx = 1

=> A\int^{2.1}_0 \frac{1-cos2(kx)}{2}dx = 1

=> A[\frac{x}{2}-\frac{sin2(kx)}{2(2)}]^{2.1}_0 = 1

=> A[1.05-\frac{sin2(k2.1)}{4}] = 1

=> A = \frac{1}{[1.05-0.25sin(4.2k)]} .